the length of a rectangle is twice the width. the are is 200yd squared. find the length and the width
Answer by edjones(8007)

L=2w
Lw=A
2w*w=200 Substitute 2w for L
2w^2=200
w^2=100
w=10yd take sqrt of each side.
L=20yd
.
Ed